Show that ๐‘ ๐‘–๐‘›๐‘–๐‘ฆ = ๐‘– ๐‘ ๐‘–๐‘›hy

1
Expert's answer
2022-03-20T06:43:32-0400

Solution in the first way. We apply the following formulas

"\sin z=\frac{e^{iz}-e^{-iz}}{2i}" and "\sinh z=\frac{e^{z}-e^{-z}}{2}". Then "\sin iy=\frac{e^{i(iy)}-e^{-i(iy)}}{2i}=\frac{e^{-y}-e^{y}}{2i}=-\frac{1}{i}\frac{e^{y}-e^{-y}}{2}=i\sinh y", as required.


Solution in the second way. We apply the following formulas

"\sin z=\sum\limits_{n=0}^{+\infty}\frac{(-1)^n}{(2n+1)!}z^{2n+1}" and "\sinh z=\sum\limits_{n=0}^{+\infty}\frac{1}{(2n+1)!}z^{2n+1}". Then

"\sin iy=\sum\limits_{n=0}^{+\infty}\frac{(-1)^n}{(2n+1)!}(iy)^{2n+1}=\sum\limits_{n=0}^{+\infty}\frac{(-1)^n}{(2n+1)!}i^{2n+1}y^{2n+1}"

"=\sum\limits_{n=0}^{+\infty}\frac{(-1)^n}{(2n+1)!}i(-1)^ny^{2n+1}=i\sum\limits_{n=0}^{+\infty}\frac{1}{(2n+1)!}y^{2n+1}=i\sinh y", as required.